OpenStudy (angelwings996):

Simplify 3 1/2 * 3 1/2

OpenStudy (anonymous):

3.5 * 3.5 = 12.25

OpenStudy (angelwings996):

\[3 \frac{ 1 }{ 2 } * 3 \frac{ 1 }{ 2 }\]

OpenStudy (angelwings996):

Is that the best it can simplify ?

OpenStudy (anonymous):

I'm pretty sure it is.

OpenStudy (anonymous):

I'm not positive though

OpenStudy (theviper):

\[3 \frac{ 1 }{ 2 } * 3 \frac{ 1 }{ 2 }=\frac{7}{2}*\frac{7}{2}=\frac{?}{?}\]

OpenStudy (angelwings996):

Well, I did it the fraction way and came up with \[\frac{ 49 }{ 4 }\]

OpenStudy (angelwings996):

@TheViper is 49/4 correct?

OpenStudy (theviper):

yes now u have to divide only :)

OpenStudy (angelwings996):

so 7/2 ?

OpenStudy (angelwings996):

Oh wait I did that wrong

OpenStudy (theviper):

no 49/4=?

OpenStudy (angelwings996):

\[1\frac{ 1 }{ 3 }\] @TheViper

OpenStudy (theviper):

no 12.25 or \(\large{12\dfrac{1}{4}}\)

OpenStudy (anonymous):

So I was right? Hahaha